Solos furniture company makes only 3 kinds of furnitures chairs, desks and tables. Each product requires some labor in the parts fabrication department,the assembly department and the shipping department.The furniture is sold through a local distributor who has estimated the maximum potential sales for each product in the next quarter.The parameters fabrication chairs 4 desk 6 tables 2 hours available 1850
Assembly chair 3,desks 5,table 7, hours available 2,400 shipping chairs 3,desks 2 table 4 hours available 1500.demand potential chair 360 desks 300,tables 100 profit chair $15 desks $24 tables $18.
What would happen in case the company decides that each product must make up at 25% of the total unit sold. What is the new level of profits?Show working.
1
Expert's answer
2021-10-02T16:38:07-0400

 4C+6d+ 2t<=1850

 3C+5d+ 7t<=2400

 3C+2d 4t<=1500

c<=360

d<=300

d<=100

c,t,h>=0

Maximize z = 15c+24d+8t

at given positions, such is the maximal function




with new data




we solve it by searching for a solution with finding the maxmal fuction